# Towards Interpretability Without Sacrifice: Faithful Dense Layer Decomposition with Mixture of Decoders

### v1

@@ 1 Introduction
Figure 1: Units of specialization for sparse layer variants: Neuron-level sparsity of existing sparse MLPs [27, 26] (center) vs layer-level sparsity (right), which the proposed Mixture of Decoders (MxD) layer enables at scale. For GPT2-124M, the dimensions are: $O=768$, $H^{*}=O\cdot 4$, $N\approx O\cdot 32$.
@@
In this paper, we advocate for moving from neuron-level to layer-level sparsity (as illustrated in Figure 1) to address this. We propose the Mixture of Decoders (MxD) layer to overcome the sparsity-accuracy trade-off through scalable, resource-efficient conditional computation. Rather than individual vectors, MxDs learn interpretable sublayers as atomic units of specialization. This faithfully mirrors the functional form of dense layer we wish to approximate, and allows MxDs to readily generalize to modern MLP variants (i.e., the Gated Linear Unit [33]).
@@
  • Across $108$ sparse layers in $4$ LLMs (with up to $3$B parameters) MxDs (i) pareto-dominate existing techniques on the sparsity-accuracy frontier yet (ii) remain competitive on $34$ sparse probing and steering tasks, validating the interpretability of the learned experts.
@@ 2 Methodology
We first recall the technical details of language models’ MLP layers and existing approaches to sparse approximations in Section 2.1. We then introduce the proposed MxD in Section 2.2, outlining the attractive rank properties it inherits in Section 2.3 and factorized implementation in Section 2.4. We conclude with extensions to modern MLP layers in Section 2.5.
@@ 2.1 Preliminaries
where $\mathbf{E}^{{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}*}}\in\mathbb{R}^{I\times H^{{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}*}}},\,\mathbf{D}^{{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}*}}\in\mathbb{R}^{H^{{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}*}}\times O}$ are the learnable ‘encoder’ and ‘decoder’ parameters respectively, and $\phi(.)$ is an activation function, often a GELU [39]. We use to denote the weights/dimensions of the pre-trained base LLM.
@@ 2.2 Mixture of Decoders
With the gate routing each token to just its top-$K$ experts, each $\mathbf{W}_{n}\in\mathbb{R}^{H\times O}$ receives a gradient signal from only a specific set of semantically similar tokens. This implicit clustering naturally leads experts to specialize in feature-specific subcomputations, while collectively covering the layer’s full functionality. MxDs in Equation 3 also directly inherit the MLP layers’ original functional form, avoiding the need to impose sparsity and non-negativity constraints on the hidden units $\mathbf{z}\in\mathbb{R}^{H}$. However, MxD decoders naively require a prohibitive $NHO$ parameters–preventing $N$ from scaling to tens of thousands of specialized components. To achieve parameter-efficiency whilst retaining layer capacity for faithful layer approximation, we parameterize MxDs’ third-order weight tensor $\boldsymbol{\mathcal{W}}\in\mathbb{R}^{N\times H\times O}$ specifically to yield full-rank expert weights, defined elementwise as:
@@
where * is the Hadamard product [34, 35], and $\mathbf{c}_{n},\,\mathbf{d}_{h}\in\mathbb{R}^{O}$ are the rows of learnable weights $\mathbf{C}\in\mathbb{R}^{N\times O},\,\mathbf{D}\in\mathbb{R}^{H\times O}$. Intuitively, $\mathbf{D}$ implements a base transformation modulated by the $N$ specialized units in $\mathbf{C}$. Additional technical motivation for this parameterization with tensor methods can be found in Section A.3. This brings MxDs’ parameter count down significantly to $O\cdot(N+H)$ from $NHO$ in Equation 3 with $N$ full decoders. One can then vary $N$ to parameter-match sparse MLP layers. We next detail how this design (i) retains expressivity in each unit for faithful layer approximation under sparsity in Section 2.3 and (ii) yields a simple forward pass in Section 2.4.
@@ 2.3 MxDs are rank-preserving
In the original LLM, the linear transformation from the hidden units to the output is constrained by the rank of the original MLP’s decoder matrix $\mathbf{D}^{{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}*}}\in\mathbb{R}^{{H}^{{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}*}}\times O}$. Under only mild technical conditions, every expert’s weight matrix in MxDs inherits the rank of $\mathbf{D}\in\mathbb{R}^{H\times O}$, thus allowing it to match that of the original MLP’s decoder, despite its parameter-efficiency:
@@
The proof is found in Section A.1, which first derives the matrix-valued expression for each expert from Equation 4 and then applies a standard rank equality. At a sparsity level of $K$, each MxD output vector is a weighted sum of $K$-many linear transformations (each with potentially full-rank weights) of the dense hidden units $\mathbf{z}$. As a result, MxDs retain layer capacity even under high sparsity. Sparse MLPs’ hidden units have only $K$ non-zero elements in contrast–each output in Equation 2 is therefore confined to a $K$-dimensional subspace of $\mathbb{R}^{O}$, potentially limiting the capacity of sparse MLPs to faithfully approximate the original mapping in the small $K$ regime desirable for interpretability (mirroring speculations by [26]). Further, whilst alternative soft linear MoEs achieve scalability through low-rankness [40], Lemma 1 states that no such rank constraints are present in MxDs. For approximating existing MLP layers where low-rank assumptions may not hold, MxDs are consequently a more suitable class of conditional layer.
@@ 2.4 Factorized forward pass
The proof is found in Section A.2. We include a notebook at https://github.com/james-oldfield/MxD/blob/main/form-equivalence.ipynb showing the equivalence in PyTorch. Further, please see Section A.5 for a discussion of how the Hadamard factorization relates to prior parameter-efficient MoEs with element-wise scaling [37].
@@
Model form ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{y}}={\mathbf{D}^{{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}*}}}^{\top}\mathbf{z}^{{\color[rgb]{0.5,0.5,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0.5}*}}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{y}}\approx\mathbf{D}^{\top}\mathbf{z}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{y}}\approx\mathbf{D}^{\top}\mathbf{z}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{y}}\approx\mathbf{D}^{\top}\mathbf{z}+\mathbf{S}^{\top}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{x}}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{y}}\approx\sum_{n}a_{n}\left(\mathbf{W}_{n}^{\top}\mathbf{z}\right)$
@@ 2.5 Extending MxDs to GLUs
where $\mathbf{a}\coloneqq\mathcal{S}\big{(}\mathbf{G}^{\top}\mathbf{x}\big{)}\in\mathbb{R}^{N}$ are the expert units, and $\mathbf{W}_{n}=\mathbf{D}\,\mathrm{diag}(\mathbf{c}_{n})\in\mathbb{R}^{H\times O}$ as before. For a technical discussion of GLUs and their relationship to MxDs, we refer readers to Section A.4–through the theoretical results developed in this paper, we show that GLU encoders themselves can be viewed as a mixture of rank-$1$ linear experts (in contrast to the rank-preserving MxDs).
@@ 3 Experiments
The experimental section in the main paper is split into two parts. Section 3.1 first demonstrates how MxDs perform significantly better on the accuracy-sparsity frontier as sparse MLP layer approximations on 4 LLMs. We then demonstrate in Section 3.2 that MxD’s features retain the same levels of specialization through sparse probing and steering evaluations. Thorough ablation studies, experiments with matrix rank, and comparisons to low rank MoEs are presented in Appendix B.
@@ 3.1 Sparse approximations of MLPs in LLMs
We train on $4$ base models: GPT2-124M [3], Pythia-410m, Pythia-1.4b [41], and Llama-3.2-3B [1] with up to $80$k experts/features. We train all sparse layers on a total of $480$M tokens of OpenWebText [42], with learning rate $41e-$ 4 and a context length of $128$, initializing the output bias as the empirical mean of the training tokens, and $\mathbf{D}$ in MxDs as the zero-matrix (following [26]). We vary $N$ in MxD layers to parameter-match Transcoders in all experiments, with parameter counts and dimensions shown in Table 2. For Llama3.2-3B, we use the Swish-GLU variant of MxD and GELU-MLP MxDs for the other three models, matching the architectures of their base encoders. Through ablation studies in Section B.6 we show that MxDs using the GELU/GLU variants are much more accurate layer approximators than the ReLU variants. Full experimental details are included in Appendix D. Whilst we do not have the computational resources to show similarly thorough experiments on even larger LLMs, we expect MxDs to scale just as well to models with tens of billions of parameters or more.
@@
Model Params $H$ $N$ Params $H$ $N$ Params $H$ $N$ Params $H$ $N$
Transcoders [27] $37.7$M $24{,}576$ $67.1$M $32{,}768$ $268.5$M $65{,}536$ $604$M $98{,}304$
Skip Transcoders [26] $38.4$M $24{,}576$ $68.2$M $32{,}768$ $272.7$M $65{,}536$ $614$M $98{,}304$
MxDs $37.7$M $3072$ $21{,}490$ $67.1$M $4096$ $28{,}658$ $268.4$M $8192$ $57{,}330$ $604$M $8202$ $86{,}015$
@@
Given the frozen weights of the MLP, we train sparse layers to minimize the normalized reconstruction loss between its output and that of the original MLP layer with objectives of the form $\mathcal{L}=\mathbb{E}_{\mathbf{x}}\left[\frac{||\mathrm{MLP}(\mathbf{x})-f(\mathbf{x})||_{2}^{2}}{||\mathrm{MLP}(\mathbf{x})||_{2}}\right]$, where $f(.)$ denotes the various learnable sparse MLP layers. To compare with recent work [26], we adopt the TopK activation function [23] for sparsity-inducing function $\mathcal{S}(.)$, removing the need for an additional sparsity penalty.
@@ 3.1.1 Results: sparsity vs faithfulness
We train an exhaustive set of $60$ sparse MLP approximations across $4$ diverse LLMs with up to $3$B parameters. We show in Figure 3 the resulting downstream base model cross-entropy loss when using the trained sparse layers in place of the original MLPs. As can be seen, not only do the proposed MxD layers outperform Transcoders [27] notably, but model performance is similarly preserved at all sparsity levels in MxD layers. With prior work finding sparse solutions to be more interpretable [17, 19], the performance gap of MxDs at small $K$ is a significant advantage. Please also see Figure 10 for results with normalized MSE, where MxDs’ reconstruction errors are up to an order of magnitude smaller. Full results on additional layers are included in Section B.3 for $48$ more trained sparse layers. Please also see Section B.1 for qualitative and quantitative results for how faithfully the sparse layers propagate to the LLMs’ output space of natural language.
The recent ‘Skip Transcoders’ (STCs) [26], introduce an additional $IO$ parameters with a skip connection $\mathbf{S}\in\mathbb{R}^{I\times O}$ mapping the input directly to the output with $\mathbf{y}\approx\mathbf{D}^{\top}\mathbf{z}+\mathbf{S}^{\top}\mathbf{x}$. STC layers thus have considerably more parameters (e.g., STCs on llama3.2-3B have $10$M more parameters than MxDs). Despite the smaller parameter counts, we find MxDs consistently outperform STCs on the sparsity-accuracy frontier, attesting to the benefits of MxDs’ model form.
@@ 3.2 Feature evaluations
The accurate reconstruction of MxD models in Section 3.1 provides some evidence that MxDs are faithfully emulating the original MLP layers’ functional mapping. However, for interpretability, we care equally about the extent to which the learned features correspond to specialized, human-interpretable concepts. We confirm that MxD’s features compete with the baselines quantitatively in two ways: through probing for known concepts in Section 3.2.1 and by steering the model using the learned features Section 3.2.2. For all experiments in this section, we use the $K=32$ models.
@@
Interestingly, we find MxDs naturally learn a ‘shared’ expert performing a common base transformation–the remaining $K-1$ active experts are thus free to dedicate their capacity to modelling features unique to individual tokens. This emergent shared/private processing complements recent trends to use shared experts by design in MoEs [43, 44, 45, 46, 47] with [43] arguing this facilitates greater specialization. Furthermore, one may view the skip connection in STCs [26] as performing an analogous role to the shared expert. With MxDs, however, all units have the same high capacity to accurately learn separate subcomputation regardless of the frequency or rarity of features.
We also observe that our trained MxDs exhibit very few ‘dead’ experts, as shown in Section C.1, with many experts contributing actively. Furthermore, initial ablations in Section C.2 show that one can train MxDs without shared experts if desired, at small performance cost. Please see qualitative results of activated tokens for particular experts in Appendix E.

### v2

@@ 1 Introduction
Figure 1: Units of specialization for sparse layer variants: Neuron-level sparsity of existing sparse MLPs [27, 26] (center) vs layer-level sparsity (right), which the proposed Mixture of Decoders (MxD) layer enables at scale. For GPT2, the dimensions are: $O=768$, $H^{*}=O\cdot 4$, $H\approx N\approx O\cdot 32$.
@@
In this paper, we advocate for moving from neuron-level to layer-level sparsity (as illustrated in Figure˜1) to address this. We propose the Mixture of Decoders (MxD) layer to overcome the sparsity-accuracy trade-off through scalable, resource-efficient conditional computation. Rather than individual vectors, MxDs learn interpretable sublayers as atomic units of specialization. This faithfully mirrors the functional form of dense layer we wish to approximate, and allows MxDs to readily generalize to modern MLP variants (i.e., the Gated Linear Unit [33]).
@@
  • Across $108$ sparse layers in $4$ LLMs (with up to 33B parameters) MxDs (i) pareto-dominate existing techniques on the sparsity-accuracy frontier yet (ii) remain competitive on $34$ sparse probing and steering tasks, validating the interpretability of the learned experts.
@@ 2 Methodology
We first recall the technical details of language models’ MLP layers and existing approaches to sparse approximations in Section˜2.1. We then introduce the proposed MxD in Section˜2.2, outlining the attractive rank properties it inherits in Section˜2.3 and factorized implementation in Section˜2.4. We conclude with extensions to modern MLP layers in Section˜2.5.
@@ 2.1 Preliminaries
[…]
@@ 2.2 Mixture of Decoders
With the gate routing each token to just its top-$K$ experts, each $\mathbf{W}_{n}\in\mathbb{R}^{H\times O}$ receives a gradient signal from only a specific set of semantically similar tokens. This implicit clustering naturally leads experts to specialize in feature-specific subcomputations, while collectively covering the layer’s full functionality. MxDs in Equation˜3 also directly inherit the MLP layers’ original functional form, avoiding the need to impose sparsity and non-negativity constraints on the hidden units $\mathbf{z}\in\mathbb{R}^{H}$. However, MxD decoders naively require a prohibitive $NHO$ parameters–preventing $N$ from scaling to tens of thousands of specialized components. To achieve parameter-efficiency whilst retaining layer capacity for faithful layer approximation, we parameterize MxDs’ third-order weight tensor $\boldsymbol{\mathcal{W}}\in\mathbb{R}^{N\times H\times O}$ specifically to yield full-rank expert weights, defined elementwise as:
@@
where * is the Hadamard product [34, 35], and $\mathbf{c}_{n},\,\mathbf{d}_{h}\in\mathbb{R}^{O}$ are the rows of learnable weights $\mathbf{C}\in\mathbb{R}^{N\times O},\,\mathbf{D}\in\mathbb{R}^{H\times O}$. Intuitively, $\mathbf{D}$ implements a base transformation modulated by the $N$ specialized units in $\mathbf{C}$. Additional technical motivation for this parameterization with tensor methods can be found in Section˜A.3. This brings MxDs’ parameter count down significantly to $O\cdot(N+H)$ from $NHO$ in Equation˜3 with $N$ full decoders. One can then vary $N$ to parameter-match sparse MLP layers. We next detail how this design (i) retains expressivity in each unit for faithful layer approximation under sparsity in Section˜2.3 and (ii) yields a simple forward pass in Section˜2.4.
@@ 2.3 MxDs are rank-preserving
[…]
@@
The proof is found in Section˜A.1, which first derives the matrix-valued expression for each expert from Equation˜4 and then applies a standard rank equality. At a sparsity level of $K$, each MxD output vector is a weighted sum of $K$-many linear transformations (each with potentially full-rank weights) of the dense hidden units $\mathbf{z}$. As a result, MxDs retain layer capacity even under high sparsity. Sparse MLPs’ hidden units have only $K$ non-zero elements in contrast–each output in Equation˜2 is therefore confined to a $K$-dimensional subspace of $\mathbb{R}^{O}$, potentially limiting the capacity of sparse MLPs to faithfully approximate the original mapping in the small $K$ regime desirable for interpretability (mirroring speculations by [26]). Further, whilst alternative soft linear MoEs achieve scalability through low-rankness [40], Lemma˜1 states that no such rank constraints are present in MxDs. For approximating existing MLP layers where low-rank assumptions may not hold, MxDs are consequently a more suitable class of conditional layer.
@@ 2.4 Factorized forward pass
The proof is found in Section˜A.2. We include a notebook at https://github.com/james-oldfield/MxD/blob/main/form-equivalence.ipynb showing the equivalence in PyTorch. Further, please see Section˜A.5 for a discussion of how the Hadamard factorization relates to prior parameter-efficient MoEs with element-wise scaling [37], and Section˜B.6 for performance/computational cost comparisons.
@@
[…]
@@ 2.5 Extending MxDs to GLUs
where $\mathbf{a}\coloneqq\mathcal{S}\big(\mathbf{G}^{\top}\mathbf{x}\big)\in\mathbb{R}^{N}$ are the expert units, and $\mathbf{W}_{n}=\mathbf{D}\,\mathrm{diag}(\mathbf{c}_{n})\in\mathbb{R}^{H\times O}$ as before. For a technical discussion of GLUs and their relationship to MxDs, we refer readers to Section˜A.4–through the theoretical results developed in this paper, we show that GLU encoders themselves can be viewed as a mixture of rank-$1$ linear experts (in contrast to the rank-preserving MxDs).
@@ 3 Experiments
The experimental section in the main paper is split into two parts. Section˜3.1 first demonstrates how MxDs perform significantly better on the accuracy-sparsity frontier as sparse MLP layer approximations on 4 LLMs. We then demonstrate in Section˜3.2 that MxD’s features retain the same levels of specialization through sparse probing and steering evaluations. Thorough ablation studies, experiments with matrix rank, and comparisons to low rank MoEs are presented in Appendix˜B.
@@ 3.1 Sparse approximations of MLPs in LLMs
We train on $4$ base models: GPT2-124M [3], Pythia-410m, Pythia-1.4b [41], and Llama-3.2-3B [1] with up to 8080k experts/features. We train all sparse layers on a total of 480480M tokens of OpenWebText [42], with learning rate 1e41e-4 and a context length of $128$, initializing the output bias as the empirical mean of the training tokens, and $\mathbf{D}$ in MxDs as the zero-matrix (following [26]). We vary $N$ in MxD layers to parameter-match Transcoders in all experiments, with parameter counts and dimensions shown in Table˜2. For Llama3.2-3B, we use the Swish-GLU variant of MxD and GELU-MLP MxDs for the other three models, matching the architectures of their base encoders. Through ablation studies in Section˜B.8 we show that MxDs using the GELU/GLU variants are much more accurate layer approximators than the ReLU variants. Full experimental details are included in Appendix˜D. Whilst we do not have the computational resources to show similarly thorough experiments on even larger models, we expect MxDs to scale just as well as sparse MLPs to models with tens of billions of parameters or more.
@@
Model Params $H$ NN Params $H$ NN Params $H$ NN Params $H$ NN
Transcoders [27] 37.737.7M 24,57624{,}576 67.167.1M 32,76832{,}768 268.5268.5M 65,53665{,}536 604604M 98,30498{,}304
Skip Transcoders [26] 38.438.4M 24,57624{,}576 68.268.2M 32,76832{,}768 272.7272.7M 65,53665{,}536 614614M 98,30498{,}304
MxDs 37.737.7M $3072$ 21,49021{,}490 67.167.1M $4096$ 28,65828{,}658 268.4268.4M $8192$ 57,33057{,}330 604604M $8202$ 86,01586{,}015
@@
[…]
@@ 3.1.1 Results: sparsity vs faithfulness
We train an exhaustive set of $60$ sparse MLP approximations across $4$ diverse LLMs with up to 33B parameters. We show in Figure˜3 the resulting downstream base model cross-entropy loss when using the trained sparse layers in place of the original MLPs. As can be seen, not only do the proposed MxD layers outperform Transcoders [27] notably, but model performance is similarly preserved at all sparsity levels in MxD layers. Please also see Figure˜9 for results with normalized MSE, where MxDs’ reconstruction errors are up to an order of magnitude smaller. Full results on additional layers are included in Section˜B.3 for $48$ more trained sparse layers.
The recent ‘Skip Transcoders’ (STCs) [26], introduce an additional $IO$ parameters with a skip connection $\mathbf{S}\in\mathbb{R}^{I\times O}$ mapping the input directly to the output with $\mathbf{y}\approx\mathbf{D}^{\top}\mathbf{z}+\mathbf{S}^{\top}\mathbf{x}$. STC layers thus have considerably more parameters (e.g., STCs on llama3.2-3B have 1010M more parameters than MxDs). Despite the smaller parameter counts, we find MxDs consistently outperform STCs on the sparsity-accuracy frontier, attesting to the benefits of MxDs’ model form.
3.1.2 Faithfulness in output space
Next, we perform experiments comparing the faithfulness of sparse layers as their computation propagates to the model output space. We sample $16$ future tokens with the base model and then measure how similar the same generations are when the target MLP layer is replaced with the sparse layers. We use $512$ text snippets from OpenWebText, and take the first 4 words of each as the initial prompts, generating $16$ future tokens after each prompt. We plot in Figures˜4(a) and 4(b) the percentage of the samples’ continuations that are identical in the original LLM and hooked LLMs up to $n$ future tokens ahead. We note that this is a rather punishing task–any small deviations quickly compound as $n$ grows. Despite this, we see that the MxDs match the future token generations far better than the baselines, exhibiting more faithfulness in model output space (as well as in latent space).
Please see qualitative examples of the first $8$ prompts and the subsequent ‘diffs’ (using Python 3’s difflib) of the generated tokens in the Appendix in Figures˜7 and 8.
(a) Pythia-410m.
(b) GPT2-124m.
Figure 4: Proportion of $512$ generated samples that contain $n$ predicted future words identical to the original model’s output when replacing the base LLM’s MLP layer with the sparse layers.
@@ 3.2 Feature evaluations
The accurate reconstruction of MxD models in Section˜3.1 provides some evidence that MxDs are faithfully emulating the original MLP layers’ functional mapping. However, for interpretability, we care equally about the extent to which the learned features correspond to specialized, human-interpretable concepts. We confirm that MxD’s features compete with the baselines quantitatively in two ways: through probing for known concepts in Section˜3.2.1 and by steering the model using the learned features Section˜3.2.2. For all experiments in this section, we use the $K=32$ models.
@@
Interestingly, we find MxDs naturally learn a ‘shared’ expert performing a common base transformation–the remaining K1K-1 active experts are thus free to dedicate their capacity to modelling features unique to individual tokens. This emergent shared/private processing complements recent trends to use shared experts by design in MoEs [43, 44, 45, 46, 47] with [43] arguing this facilitates greater specialization. Furthermore, one may view the skip connection in STCs [26] as performing an analogous role to the shared expert. With MxDs, however, all units have the same high capacity to accurately learn separate subcomputation regardless of the frequency or rarity of features.
We also observe that our trained MxDs exhibit very few ‘dead’ experts, as shown in Section˜C.1, with many experts contributing actively. Furthermore, initial ablations in Section˜C.2 show that one can train MxDs without shared experts if desired, at small performance cost. Please see qualitative results of activated tokens for particular experts in Appendix˜E.